A piece of steel is in the shape of a triangle. One side is 6 inches long. The lengths of the two other sides are equal and each is less than 6 inches long. The lengths of the unknown sides of the triangle are whole numbers of inches. What are all the possible lengths, in inches, of the unknown sides of the triangle? Explain the process you used to find all the possible lengths.

Answer by math_helper(2461)

To be an integer number of inches long, and less than 6, the possible lengths are:
This follows from the triangle inequality, which states that the sum of any two lengths of a triangle must be greater than the length of the third side:


4+4 = 8 > 6
4+6 = 10 > 4


5+5 = 10 > 6
6+5 = 11 > 5